Two infinitely parallel plates have charges 1.6 microcoulomb per meter square and 2 microcoulomb per meter square if they are arranged close to each other what is the electric field
( A ) in the region outside the plates
(B) in between the page
1
Expert's answer
2020-06-05T11:32:57-0400

The fields from the first and the second plate:


"E_1=\frac{\sigma_1}{2\epsilon_0},\space E_2=\frac{\sigma_2}{2\epsilon_0}."

Therefore, the field outside the plates will be


"E_o=E_2-E_1=\frac{\sigma_2-\sigma_2}{2\epsilon_0}=22.6\text{ kV/m},"

the field between the plates is

"E_i=E_2+E_1=\frac{\sigma_2+\sigma_2}{2\epsilon_0}=203\text{ kV/m}."